OpenStudy (ashleyvaldes501):

Solve x2 + 8x − 3 = 0 using the completing-the-square method.

OpenStudy (welshfella):

divide coefficient of x by 2 8 / 2 = 4 Now we write x^2 + 8x in the 'complete the square form' :- x^2 + 8x = (x + 4)^2 - 16 Can you see why its -16?

OpenStudy (welshfella):

Yes? , No?

OpenStudy (ashleyvaldes501):

yes thank u!

OpenStudy (welshfella):

ok so (x + 2)^2 - 16 - 3 = 0 (x+ 4)^2 = 19 now take the square root of both sides

OpenStudy (welshfella):

x + 4 = +/- sqrt19 now subtract 4 from both sides and you have the solution